A student measures the molar solubility of barium sulfite in a water solution to be 8.64E-4 M.  
Based on her data, the solubility product constant for this compound is ?
Answer #1


At equilibrium:
BaSO3     <---->     Ba2+     +          SO32-  

                     s                  s       


Ksp = [Ba2+][SO32-]
Ksp = (s)*(s)
Ksp = 1(s)^2
Ksp = 1(8.64*10^-4)^2
Ksp = 7.465*10^-7
Answer: 7.46*10^-7